$A$ box of oranges is inspected by examining three randomly selected oranges drawn without replacement. If all the three oranges are good,the box is approved for sale; otherwise,it is rejected. Find the probability that a box containing $15$ oranges,out of which $12$ are good and $3$ are bad,will be approved for sale.

  • A
    $\frac{44}{91}$
  • B
    $\frac{45}{91}$
  • C
    $\frac{46}{91}$
  • D
    $\frac{47}{91}$

Explore More

Similar Questions

$A$ bag $B$ contains $4$ white balls and $2$ black balls. Another bag $C$ contains $3$ white balls and $5$ black balls. If one ball is drawn randomly from each bag, then the probability that the two balls drawn are both white is

If $P(A_1 \cup A_2) = 1 - P(A_1^c) P(A_2^c)$,where $c$ stands for complement,then the events $A_1$ and $A_2$ are

Two cards are drawn at random and without replacement from a pack of $52$ playing cards. Find the probability that both the cards are black.

An urn contains $10$ black and $5$ white balls. Two balls are drawn from the urn one after the other without replacement. What is the probability that both drawn balls are black?

$A$ and $B$ are two independent events such that $P(A) = \frac{1}{2}$ and $P(B) = \frac{1}{3}$. Then $P$ (neither $A$ nor $B$) is equal to

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo